OpenStudy (anonymous):

verify cos(180 + x)= -cosx

OpenStudy (solomonzelman):

\(\large\color{blue}{ \bf cos(180+x)=cos(180)cos(x)-sin(180)sin(x)=}\) \(\large\color{purple}{ \bf cos(180)=-1}\) \(\large\color{purple}{ \bf sin(180)=0}\) \(\large\color{black}{ \bf hence,~~you~~get~}\) \(\large\color{blue}{ \bf cos(180+x)=cos(180)cos(x)-sin(180)sin(x)}\) \(\large\color{blue}{ \bf =(-1)cos(x)-(0)sin(x)=-cos(x)}\)